# Revisiting Who’s Harry Potter: Towards Targeted Unlearning from a Causal Intervention Perspective

### v1

@@ Appendix A Construction of WPU
Appendix B Evaluation Metrics
@@
$\bullet$ GPT quality score: Given the question and generated response, we use GPT-4 to evaluate the quality of the response, with scores ranging from $\{1,2,3\}$, where $3$ denotes the response is fluent, relevant, and appropriate, regardless of its correctness. This score is computed for both forget QA and hard-retain QA. Particularly, on forget QA, we replace the person’s full name with the last name in both the question and response before feeding them to GPT-4, since we observe that GPT-4 tends to assign low scores to responses that do not match the factual information, even if the response is fluent and relevant. Using the person’s last name instead of full name effectively prevents GPT-4 from focusing on the correctness of the response, while still being able to detect irrelevant responses, e.g., when the question asks about Wilhelm Wattenbach but the model answers with other persons (see examples in Appendix F). The prompts we use on forget QA and hard-retain QA are shown in Figures A8 and A9 respectively.
@@
Appendix C Implementation Details
We now describe the implementation details for baselines and out method. Tables A2 and A3 show the training hyper-parameters for all methods. We evaluate on Llama2-7b-chat Touvron et al. (2023) on WPU and the fine-tuned model provided by Maini et al. (2024) on TOFU. All experiments are run on two NVIDIA A6000 GPUs. The average training time for each unlearned model is less than 10 minutes.
@@ Appendix C Implementation Details
Learning rate $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
Table A2: Training hyper-parameters on WPU. GA and NPO use learning rate of $4\times 10^{-5}$, all other methods use learning rate of $1\times 10^{-5}$.
@@
C.1 Implementation Details on WPU
Baselines. For GA and NPO, we use the official implementation in Maini et al. (2024) and Zhang et al. (2024). The retain documents contain Wiki pages of 100 persons that do not overlap with any test data. For Prompt, we use the same instruction in Thaker et al. (2024), with a few modifications made for the targeted unlearning task. Figure A11 shows the detailed prompt we use. For Prompt-distill, we construct the teacher distribution and train the student model on two sets of QA pairs. The first set contains questions about the unlearning target, and the student LLM should learn to refuse these questions. Specifically, we evaluate the output distribution of Prompt on its own generated responses and set it as the teacher distribution. Note that these teacher responses are mostly like ‘I don’t known this person’. The student model is then trained to mimic this distribution, without the prepended unlearning prompt. We create additional questions about the unlearning targets for training, and make sure they do not overlap with the questions in the test data. The second set contains normal questions that the student LLM should answer correctly. We obtain the teacher distribution from the original LLM (without the unlearning prompt) on a set of questions unrelated to the unlearning targets. We further filter the teacher responses and only keep the correct ones. For DI, we use the official implementation in Dong et al. (2024) and reduce the logit of the original token by 10. For WHP, we re-implement it based on our best understanding of the method Eldan and Russinovich (2023). Particularly, we only implement the name change algorithm, without the reinforcement bootstrapping, to keep consistency with our framework. Additionally, Eldan and Russinovich (2023) shows that the name change algorithm is the major design contributing to unlearning.
@@ C.1 Implementation Details on WPU
C.2 Implementation Details on TOFU
Baselines. The baseline implementations are similar to Appendix C.1. For GA and NPO, we use the original retain data in TOFU for the regularization term. For Prompt, we prepend the unlearning prompt to the model and follow Maini et al. (2024) to measure forget quality and model utility. For Prompt-distill, we observe that many responses from Prompt still contain the correct information about the unlearning target, since the model is overfitting on the data. We thus filter the responses from Prompt to only keep those having a ROUGE score lower than 0.4 for training.
@@ C.2 Implementation Details on TOFU
Our method. We consider both authors and their books as the unlearning targets and change their names in the input. We use the same list of person names as in Table A4 for replacement. There are two designs that are different from Appendix C.1. First, unlike persons, a book title can indicate its content, e.g., Shale Stories suggests it is about shale. Since this knowledge should not be forgotten, we use GPT-4 to generate alternative titles with similar meanings for replacement, e.g., Slate Tales, so that the teacher retains the knowledge that can be inferred from the title, but nothing else. Second, as discussed in section 4.3, we replace the prefix of a person’s or book’s name when predicting the next token in the name, e.g., predicting Stories given Slate. To achieve this, when generating book names for replacement, we ask GPT-4 to generate names with similar syntactic structure with the original name (e.g., having common words at their original positions). Table A5 shows some examples of the book names we use for replacement.
Appendix D Additional Results on WPU
@@
Appendix E Additional Results on TOFU
@@
Appendix F Additional Ablation Study

### v2

@@ Appendix B Construction of WPU
Appendix C Evaluation Metrics
@@
$\bullet$ GPT quality score: Given the question and generated response, we use GPT-4 to evaluate the quality of the response, with scores ranging from $\{1,2,3\}$, where $3$ denotes the response is fluent, relevant, and appropriate, regardless of its correctness. This score is computed for both forget QA and hard-retain QA. Particularly, on forget QA, we replace the person’s full name with the last name in both the question and response before feeding them to GPT-4, since we observe that GPT-4 tends to assign low scores to responses that do not match the factual information, even if the response is fluent and relevant. Using the person’s last name instead of full name effectively prevents GPT-4 from focusing on the correctness of the response, while still being able to detect irrelevant responses, e.g., when the question asks about Wilhelm Wattenbach but the model answers with other persons (see examples in Appendix G). The prompts we use on forget QA and hard-retain QA are shown in Figures A11 and A12 respectively.
@@
Appendix D Implementation Details
We now describe the implementation details for baselines and our method. Tables A2 and A3 show the training hyper-parameters for all methods. We evaluate on Llama2-7b-chat Touvron et al. (2023) on WPU and the fine-tuned model provided by Maini et al. (2024) on TOFU. All experiments are run on two NVIDIA A6000 GPUs. The average training time for each unlearned model is less than 10 minutes.
@@ Appendix D Implementation Details
Learning rate $51e-5,2e-5,3e-$ 5
Table A2: Training hyper-parameters on WPU. For all methods, we report the performance of the best learning rate among the three.
@@
D.1 Implementation Details on WPU
Baselines. For GA and NPO, we use the official implementation in Maini et al. (2024) and Zhang et al. (2024). The retain documents contain Wiki pages of 100 persons that do not overlap with any test data. For Prompt, we use the same instruction in Thaker et al. (2024), with a few modifications made for the targeted unlearning task. Figure A14 shows the detailed prompt we use. For Prompt-distill, we construct the teacher distribution and train the student model on two sets of QA pairs. The first set contains questions about the unlearning target, and the student LLM should learn to refuse these questions. Specifically, we evaluate the output distribution of Prompt on its own generated responses and set it as the teacher distribution. Note that these teacher responses are mostly like ‘I don’t know this person’. The student model is then trained to mimic this distribution, without the prepended unlearning prompt. We create additional questions about the unlearning targets for training, and make sure they do not overlap with the questions in the test data. The second set contains normal questions that the student LLM should answer correctly. We obtain the teacher distribution from the original LLM (without the unlearning prompt) on a set of questions unrelated to the unlearning targets. We further filter the teacher responses and only keep the correct ones. For DI, we use the official implementation in Dong et al. (2024) and reduce the logit of the original token by 10. For WHP, we re-implement it based on our best understanding of the method Eldan and Russinovich (2023). Particularly, we only implement the name change algorithm, without the reinforcement bootstrapping, to keep consistency with our framework. Additionally, Eldan and Russinovich (2023) shows that the name change algorithm is the major design contributing to unlearning.
@@ D.1 Implementation Details on WPU
D.2 Implementation Details on TOFU
Baselines. The baseline implementations are similar to Appendix D.1. For GA and NPO, we use the original retain data in TOFU for the regularization term. For Prompt, we prepend the unlearning prompt to the model and follow Maini et al. (2024) to measure forget quality and model utility. For Prompt-distill, we observe that many responses from Prompt still contain the correct information about the unlearning target, since the model is overfitting on the data. We thus filter the responses from Prompt to only keep those having a ROUGE score lower than 0.4 for training.
@@ D.2 Implementation Details on TOFU
Our method. We consider both authors and their books as the unlearning targets and change their names in the input. We use the same list of person names as in Table A4 for replacement. There are two designs that are different from Appendix D.1. First, unlike persons, a book title can indicate its content, e.g., Shale Stories suggests it is about shale. Since this knowledge should not be forgotten, we use GPT-4 to generate alternative titles with similar meanings for replacement, e.g., Slate Tales, so that the teacher retains the knowledge that can be inferred from the title, but nothing else. Second, as discussed in section 4.3, we replace the prefix of a person’s or book’s name when predicting the next token in the name, e.g., predicting Stories given Slate. To achieve this, when generating book names for replacement, we ask GPT-4 to generate names with a similar syntactic structure to the original name (e.g., having common words at their original positions). Table A5 shows some examples of the book names we use for replacement.
Appendix E Additional Results on WPU
@@
E.1 Mitigation Mechanism for Reversing Causal Relations
Based on the promising performance of Ours non-factual, we can design a potential mitigation for our algorithm when the causal relation is flipped, i.e., $Y$ points to $\bm{X}$ instead of the other way around. Specifically, we can convert the unlearning document into Wikipedia style (not necessarily contain factual information), and since Wikipedia text mostly follows our causal graph in Figure 3, we can apply our algorithm to converted text.
E.2 Evaluation with Llama
Method GPT privacy score Llama privacy score GPT quality score Llama quality score GPT rejection rate Llama rejection rate
GA 0.90 0.88 0.01 0.18 0.69 0.65
NPO 0.93 0.85 0.16 0.55 0.32 0.29
Prompt 0.88 0.92 0.85 0.63 0.88 0.92
Prompt-distill 0.88 0.90 0.72 0.67 0.85 0.87
DI 0.87 0.90 0.57 0.70 0.60 0.62
WHP 0.75 0.80 0.67 0.54 0.49 0.54
WHP+ (Ours-1) 0.93 0.90 0.63 0.67 0.82 0.82
Ours 0.90 0.93 0.86 0.78 0.84 0.84
Table A7: Comparison of GPT-4 and Llama-3 evaluation results.
To ensure there is no systematic bias from the use of GPT-4 in both data generation and evaluation, we repeat the GPT-4 evaluations in Appendix C using Llama3.1-70b-instruct Llama Team (2024). Results in Table A7 show that the two models provide consistent evaluations.
E.3 Comparison with RLHF Baseline
Figure A2: Comparison with RLHF baselines on WPU.
We compare with the RLHF baseline Yao et al. (2024b) on WPU. Specifically, the baseline consists of an SFT stage and a DPO stage Rafailov et al. (2023). For SFT, we train the model to output “I don’t know” responses on queries about unlearning targets, and output standard responses (Lllama’s original response) on retain data. For DPO, on retain data, we set the standard response as the chosen one and “I don’t know” response as the rejected one. On forget data, we use the opposite direction. As can be observed in Figure A2, the RLHF baseline achieves high unlearning efficacy and hallucination avoidance. However, similar to Prompt-distill, the model utility is compromised, where many regular questions are mistakenly rejected.
E.4 Additional Results on Llama-3
Figure A3: Performance on WPU using Llama-3.
We further evaluate our method and two most competitive baselines on Llama-3-8b-instruct Llama Team (2024). Results in Figure A3 show similar trends to results on Llama-2. Specifically, without access to any retain data or explicitly optimizing for fewer hallucinations, our method achieves competitive performance on all five criteria, whereas baselines suffer on some criteria, such as the drop of model utility for Prompt-distill.
E.5 Generalization to Other Languages and Entity Names
Method Origianl Spanish French MSJ Embedding
GA 88.30 87.00 81.12 83.54 87.28
NPO 79.93 81.25 79.48 74.27 60.07
Prompt 82.94 90.11 84.93 5.00 29.65
Prompt-distill 84.15 84.57 88.42 89.53 73.83
DI 83.38 77.91 82.07 76.12 62.59
WHP 75.89 76.37 70.36 71.24 74.78
WHP+(Ours-1) 87.07 83.71 88.06 90.46 73.04
Ours 86.42 83.91 87.01 86.72 75.80
Table A8: Unlearning efficacy (higher is better) when models are evaluated in different languages and jailbreaking attacks.
Method Origianl Alias
GA 90.13 89.72
NPO 77.29 77.63
Prompt-distill 85.63 82.83
DI 85.67 85.28
WHP 72.16 86.19
WHP+(Ours-1) 85.95 89.06
Ours 84.13 88.90
Table A9: Unlearning efficacy (higher is better) when models are evaluated on question with aliases of unlearning targets.
In addition to the above evaluations, we also test the unlearned models’ generalizability to different languages and aliases of unlearning targets during inference.
First, We evaluate the unlearned models when forgetting queries are presented in Spanish or French. Table A8 shows the unlearning efficacy (higher is better) for the original (in English) and translated queries on WPU-2 person setting. For reference, we also include the unlearning efficacy under the two jailbreaking attacks we considered in Figure 4, i.e., many-shot jailbreaking (MSJ) and embedding space attack (Embedding). The best performance in each column is highlighted in bold (except GA because its responses are gibberish). As can be observed, the performance of most methods can generalize to different languages. In addition, the two attacks in Figure 4 are stronger and lead to larger performance degradation, especially for training-free methods such as Prompt.
Second, we evaluate models’ generalizability to aliases of unlearning targets. Specifically, on WPU, we prompt GPT-4 to generate aliases for the unlearning targets, which we manually verify. This process identifies a subset of 9 persons with an alias, such as José Batlle y Ordóñez having the alias Pepe Batlle. While some aliases still appear in the unlearning documents, their occurrence is not frequent. We re-evaluate the performance on this subset by replacing the unlearning targets’ names with their aliases in the questions. Table A9 shows the unlearning efficacy (higher is better) given the original question and question with alias. The results suggest that most methods are robust to the variation of entity names.
E.6 Tradeoff between Five Criteria
Figure A4: Correlation matrix between five criteria on WPU.
To investigate the tradeoff between various metrics, we show the correlation matrix between the five criteria in Figure A4. Specifically, we use the results from Figure A1, where performance of all methods under all learning rates are collected to calculate the correlation between each pair of criteria.
There are three observations. First, we notice that the main tradeoff is between model utility and unlearning efficacy (similarly for adversarial robustness), where improving unlearning efficacy generally compromises model utility for all methods. This is consistent with observations in existing works Maini et al. (2024). Second, we observe a moderate negative correlation between unlearning efficacy and response quality. This is due to the fact that many unlearning methods decrease the probability of the ground-truth tokens (e.g., GA and NPO), thus more thorough unlearning leads to issues such as model degeneration. Third, we observe a positive correlation between unlearning efficacy and hallucination avoidance, since rejecting questions about unlearning targets naturally leads to less leakage of factual information.
@@ E.6 Tradeoff between Five Criteria
Appendix F Additional Results on TOFU
@@
Appendix G Additional Ablation Study